OpenStudy (anonymous):

Write the equation in slope- intercept form. Then identify the slope and the y intercept. 3x+2y=-12

OpenStudy (anonymous):

roseanne try to convert into y = mx +c now m is constant

OpenStudy (anonymous):

actually sorry m is slope and c is constant

OpenStudy (anonymous):

3x + 2y = -12 2y = -3x -12 y = -3/2 x - 6

OpenStudy (anonymous):

so its the last one. man I need to practice these big TIME

OpenStudy (anonymous):

yes he used the steps to show you how to do it rose